### v1

# FGBERT: Function-Driven Pre-trained Gene Language Model for Metagenomics

Abstract
Metagenomic data, comprising mixed multi-species genomes, are prevalent in diverse environments like oceans and soils, significantly impacting human health and ecological functions. However, current research relies on K-mer representations, limiting the capture of structurally relevant gene contexts. To address these limitations and further our understanding of complex relationships between metagenomic sequences and their functions, we introduce a protein-based gene representation as a context-aware and structure-relevant tokenizer. Our approach includes Masked Gene Modeling (MGM) for gene group-level pre-training, providing insights into inter-gene contextual information, and Triple Enhanced Metagenomic Contrastive Learning (TEM-CL) for gene-level pre-training to model gene sequence-function relationships. MGM and TEM-CL constitute our novel metagenomic language model FGBERT, pre-trained on 100 million metagenomic sequences. We demonstrate the superiority of our proposed FGBERT on eight datasets.
1 Introduction
Metagenomics, the study of mixed genomes of microbial communities in the environment (e.g. gut microbiomes or soil ecosystems) [37, 39], has revealed the critical role in fundamental biological processes like enzyme synthesis, gene expression regulation, and immune function [48, 36]. In light of this, deciphering the complex sequences of multiple species in metagenomics is vital for unraveling life’s mechanisms and advancing biotechnology [4, 33, 34].
Figure 1: Two types of complex relationships between gene sequences and functions in metagenomics. One-to-Many means that the same gene may display different functions based on the genomic context; for example, ATP synthase works differently in plants, heterotrophic bacteria, and humans. Many-to-One shows that multiple genes may perform the same function; for instance, different genes from different bacteria, e.g., Cpf1, Cas1, etc. produce the same resistance function within the immune system CRISPR.
In metagenomic data analysis, various computational methods have emerged. Traditional alignment-based methods rely on homology search against reference genome databases [53]. Machine learning-based methods [20, 31, 42] generally take K-mer as the basic processing unit and adopt its frequency or embedding as the input feature, which proved to be efficient alternatives to traditional methods. Recently, Transformer [57] has been widely used for modeling sequence data due to its ability to capture long-range dependencies. Several pre-trained genomic models [63, 10] utilize Transformer architectures to effectively model complex DNA contexts. However, these models, primarily designed for single-species analysis, face challenges in mixed multi-species metagenomic data. MetaTransformer [56] and ViBE [18] use K-mer embeddings, cater to specific metagenomic applications, with the former using Transformer for read classification and the latter pre-trained BERT to predict virus classes.
Yet existing genome or metagenomic pre-trained models still face many challenges when dealing with metagenomic data. First, how to encode a gene sequence with biological meaning? Typical pre-processing of genomic data often relies on K-mer to represent a local feature of a gene sequence. However, this approach, with its fixed and limited ‘vocabulary’ of K-mer combinations [63], falls short in capturing informative and global gene representations. Secondly, how to clarify the complicated relationships between gene sequences and their functions in metagenomic analysis? This challenge can be categorized as One-to-Many and Many-to-One problems, as opposed to single-species genome analyses (e.g., identifying regulatory elements in a particular genome). One-to-Many problem indicates that a single gene can exhibit various functions in different genomic contexts, underscoring the significance of inter-gene interactions in function regulation [59]. Figure 1a demonstrates that the ATP synthase exhibits specific functionalities in diverse organisms such as bacteria, plants, and humans. Conversely, Many-to-One problem implies that different genes can share the same function, emphasizing expression commonality [3]. As shown in Figure 1b, CRISPR [23] is an immune mechanism where different proteins such as Cpf1, Cas1, and Cas13, etc. serve this function.
To address these challenges, we propose FGBERT, a novel metagenomic pre-trained model equipped with the ability to encode contextually-aware and functionally relevant representations of metagenomic sequences. We propose a protein-based gene representation as a context-aware tokenizer to encode each gene sequence, which allows for a flexible token vocabulary for longer metagenomic sequences. This strategy leverages the inherent protein functional and structural information encoded in metagenomic data [48]; and, due to the redundancy of codon-to-amino acid mappings [12], since we adopt the embedding of protein sequences, the function remains unchanged despite potential mutations. With the new tokenizer, we propose a gene group-level pre-training task, named Masked Gene modeling (MGM), to mask some genes randomly and then pre-train a Roberta-like model to predict them, enhancing the understanding of inter-/intra-gene contextual relationships and how the same gene act in different genome contexts in One-to-Many scenarios. Furthermore, we propose Triplet Enhanced Metagenomic Contrastive Learning (TEM-CL) to identify genes with similar functions together with data augmentation and negative sampling module, boosting the model’s ability to recognize commonalities among gene functions in Many-to-One scenarios. MGM and TEM-CL constitute a joint pre-training model, which advances the co-representation learning of metagenomic gene sequences and functions.
Contributions. To the best of our knowledge, this is the first metagenomic pre-trained model encoding context-aware and function-relevant representations of metagenomic sequences. To summarize: (1) We introduce a new idea of protein-based gene representations to learn biologically meaningful tokenization of long sequences. (2) Hinged on this idea, we propose MGM to model inter-gene relationships and TEM-CL to learn complex relationships between gene sequences and functions. (3) We conduct extensive experiments on four levels of downstream tasks. Our proposed FGBERT obtains state-of-the-art performance in comparison to a comprehensive suite of baselines.
Figure 2: Overview of FGBERT. A metagenomic sequence $\mathcal{X}$ is converted into ordered protein-based gene representations $\mathcal{G}$ via a Context-Aware Tokenizer. Next, we pre-train a Gene Encoder with $\mathcal{L}_{\text{MGM}}$, 15% of these tokens are masked to predict labels $\mathcal{Y}$. Meanwhile, we introduce $\mathcal{L}_{\text{Tri}}$ to distinguish gene sequences. The data augmentation and negative sampling modules generate positive samples $\mathcal{G}^{+}$ and negative samples $\mathcal{G}^{-}$, respectively. Finally, after fine-tuning, FGBERT can handle various downstream tasks.
2 Related Works
Research on Metagenomics Traditional alignment-based methods like MetaPhlAn5 [53], aim to match similarities between query sequences and known reference genomes and are common for taxonomic profiling. Advancements in deep learning have led to new methods like CNN-MGP [2] and DeepVirFinder [49], which use CNNs for gene and viral classifications with one-hot encoding. K-mer tokenization [14], employed in approaches like MDL4Microbiome [30], is a standard for DNA sequence characterization. Additionally, Virtifier [41] maps a nucleotide sequence using a codon dictionary, combined with LSTM to predict viral genes. DeepMicrobes [31] employs a self-attention mechanism, while DeepTE [58] uses K-mer inputs with CNNs for element classification, and Genomic-nlp [42] applies word2vec for gene function analysis. MetaTransformer [56] uses K-mer embedding for species classification with Transformer. For pre-training models, LookingGlass [20] uses a three-layer LSTM model for functional prediction in short DNA reads. ViBE [18] employs a K-mer token-based BERT model pre-trained with Masked Language Modeling (MLM) for virus identification.
Pre-Training on Genomics The BERT model, effective in DNA sequence characterization, is limited by the Transformer architecture’s computational burden. LOGO [60] addresses this by cutting off long sequences into 1-2kb sub-sequences. Enformer [6] combines extended convolution with Transformers for long human genomic data. GenSLMs [64] introduce hierarchical language models for whole-genome modeling. DNABERT [24], the first pre-trained model on the human genome that focuses on extracting efficient feature representations from gene sequences. DNABERT2 [63], its successor, uses Byte Pair Encoding on multi-species genomic data. NT [10] is trained on nucleotide sequences from human and other species genomes and evaluated on 18 genome prediction tasks. HyenaDNA [43] presents a long-range genome sequence model based on single-nucleotide polymorphisms on the human reference genome.
3 Methods
In this section, we detail the proposed pre-training framework FGBERT, which contains MGM and TEM-CL illustrated in Figure 2.
Notation Given a metagenomic long sequence $\mathcal{X}$, after ultrasonic fragmentation and assembly techniques, N genes $\left\{{x_{i}}\right\}_{i=1}^{N}$ are obtained, which is tokenized to $g_{i}\in\mathbb{R}^{d}$ together with its corresponding label $y_{i}\in\mathbb{R}^{100}$, where $d$ is the token dimension. After that, we concatenate ${\left\{g_{i}\right\}}_{i=1}^{n}$ into a gene group $\mathcal{G}$. During the pre-training stage, each gene group $\mathcal{G}$ is passed to the context-aware genome language encoder $\mathcal{F}(\cdot)$ to obtain knowledge representations $z_{i}$ for the input token $g_{i}$, where $z_{i}=\mathcal{F}(g_{i})$. Then we apply a fully connected layer to obtain $\hat{z}_{i}$ from $z_{i}$, where $\hat{z}_{i}=\mathcal{H}(z_{i})$, $\mathcal{H(\cdot)}$ is a multi-classification head. Additionally, we integrate contrastive learning. For gene $x_{i}$, we introduce a data augmentation module to construct positive samples $x_{j(i)}$ and employ hard negative sampling strategy mining for negative sample construction $x_{k(i)}$.
3.1 Masked Gene Modeling (MGM)
Context-aware Tokenizer To address the first challenge, we propose the protein-based gene representation as a context-aware tokenizer, utilizing the ESM-2 [32] with 15 billion parameters. Naturally, this could be replaced by any Protein Language Model (PLM). This process starts with transforming the DNA sequence $x_{i}$ from metagenomic sequences $\mathcal{X}$ into the protein sequence by ENA software [17]. The AA sequence is then transformed into a 1280D normalized ESM-2 representation. An additional dimension is appended to encode gene orientation, resulting in 1281D gene representation $g_{i}$. These representations are concatenated to form gene groups $\mathcal{G}$ as shown in Figure 3. The conversion from DNA to AA sequences is motivated by the equivalence of certain degenerate DNA codes (i.e. non-‘ATCG’ symbols in $\mathcal{X}$) when translated into AA, underscoring the rationale for transitioning from DNA to AA sequences for clarity and consistency in representation. The gene-protein relationship inherited in metagenomic data itself [48] could be revealed by ESM-2’s comprehensive training data and protein structural and functional information, and thus ESM-2 representation preserves important intra-gene contextually-aware information. This approach bypasses traditional vocabulary limits and avoids raw sequence processing, this provides the flexibility to analyze longer metagenomic sequences.
Figure 3: Framework of our proposed Context-Aware Tokenizer.
Masked Gene Modeling With this new tokenizer, we propose MGM to pre-train the Gene Encoder. During pre-training, each token (gene) is masked with a 15% probability and predicted based on its genome context. The loss $\mathcal{L}_{MLM}$ is formulated as a cross-entropy loss:
$\mathcal{L}_{\text{MLM}}=\mathbb{E}_{g\sim\mathcal{G}}\mathbb{E}_{M}\sum_{i\in M}-\log p(g_{i}|{\mathcal{G}_{/M}})$ (1)
for any given gene representation $g$, a random set of indices, $M$ means the masked gene’s index set, the unmasked gene embeddings ${\mathcal{G}_{/M}}$ as the context.
In this way, BERT can learn inter-gene relationships within the genome, contributing to our understanding of gene functionality in diverse genomic contexts in One-to-Many scenarios. BERT implicitly learns intra-gene information encapsulated within the PLM representations, further enriching the model’s comprehension of gene characteristics.
What’s more, acknowledging gene polymorphism [47, 62], MGM emphasizes identifying multiple genes that may coexist at a single genomic site, denoted as $\hat{z}_{i}=[\hat{z}_{1},\hat{z}_{2},\hat{z}_{3},\hat{z}_{4}]$. This necessitates its ability to predict not the individual genes but also various gene combinations within the same site. Hence, two loss functions are combined into $\mathcal{L}_{\text{MGM}}$:
$\mathcal{L}_{\text{MGM}}=\frac{1}{N}\textstyle\sum\limits_{i=1}^{N}{(1-\frac{y_{i}^{T}\hat{z}_{i}}{\left\|y_{i}\right\|\cdot\left\|\hat{z}_{i}\right\|})}^{\gamma}+\\ \frac{\alpha}{N}\textstyle\sum\limits_{i=1}^{N}||\tilde{z}_{i}-\tilde{y}_{i}||_{2}^{2}$ (2)
where $\gamma$ is a reconstruction loss with the scaled cosine error, and $\alpha$ is a weighting factor to balance the importance of the two loss functions. The first item is Feature Reconstruction Loss (FRL), which measures the distance between the model prediction $\hat{z}_{i}$ and its corresponding label $y_{i}$. The second item, Probability Prediction Loss (PPL), assesses the discrepancy between the embedding probability $\tilde{z}_{i}=\frac{e^{\hat{z}_{i}}}{\sum_{j=1}^{C}e^{\hat{z}_{j}}}$ and the true category probability $\tilde{y}_{i}=\frac{e^{y_{i}}}{\sum_{j=1}^{C}e^{y_{j}}}$ processed by the softmax function with C denoting the number of gene combinations and is set to 4.
3.2 Triplet Enhanced Metagenomic Contrastive Learning (TEM-CL)
Although MGM can learn the gene-protein contextual relationship within metagenomic data, we find its ability to identify sequences with same function to be lacking, which is common in Many-to-One scenarios. During the Enzyme Commission (EC) number annotation of partial metagenomic sequences, we observe that sequences with the same EC number are close to each other in feature space and, vice versa, farther apart. This finding motivates us to adopt a contrastive learning approach to capture the functional relationships between gene classes, enabling different genes with similar functions to cluster together and further optimize model training.
The common contrastive learning objective is to learn an embedding function $\mathcal{F}$ such that the distance between positive pairs is less than that of negative pairs: $d(\mathcal{F}(x_{a}),\mathcal{F}(x_{p}))<d(\mathcal{F}(x_{a}),\mathcal{F}(x_{n}))$, where $d(\cdot,\cdot)$ is a distance function (e.g., Euclidean distance) defined on the embedding space. We adopt SupCon-Hard loss [29] to thus consider multiple positive and negative samples for each anchor, which encourages the model to mine from difficult samples and enhances the robustness and generalization of the model. Additionally, the data augmentation and negative sampling modules are also included to create positive and negative samples and improve the model’s capacity to recognize commonalities among gene classes.
Triplet Sampling How to sample the triplets is crucial to learning a well-organized embedding space. For each gene group $\mathcal{G}$, as an anchor gene $x_{i}$ within a gene batch $I$, a mutation strategy is proposed to augment orphan sequences (i.e., functions associated with individual sequences) to generate a large number of pairs of positive samples $x_{j(i)}\in\mathcal{G}_{i}^{+}$ where $\mathcal{G}_{i}^{+}$ is the positive samples set of the anchor $x_{i}$. Concretely, 10 random mutations are performed on each gene sequence, with mutation ratios randomly generated according to a standard normal distribution, and the number of mutations is calculated based on the length of the sequence. This process aims to generate new sequences that are functionally similar to the original sequence but sequentially different, thus providing additional training data for the model and improving the predictive power and accuracy of orphan EC numbers.
Hard Negatives Sampling Previous studies [19] have shown that another key to successful contrastive learning is the balance between the triviality and the hardness of the sampled triplets. For the negative sample pair $x_{k(i)}\in\mathcal{G}_{i}^{-}$ where $\mathcal{G}_{i}^{-}$ is the negative samples set of the anchor $x_{i}$.
To begin, we calculate cluster centers for each EC number sequence by averaging the embedded representations of all sequences within the same functional group. These cluster centers represent the centroids of each functional group. Subsequently, we compute Euclidean distances $d(\cdot)$ based on these centers. When selecting negative samples, we choose samples that are similar to the anchor in latent space but belong to different clusters, rather than randomly selecting them, to increase the challenge of learning. The triplet loss $\mathcal{L}_{\text{Tri}}(x_{i},\{x_{j(i)}\}_{j=1}^{N_{j}},\{x_{k(i)}\}_{k=1}^{N_{k}})$ is defined as follows:
$\mathcal{L}_{\text{Tri}}=-\textstyle\sum\limits_{i\in I}\log(\frac{1}{|\mathcal{G}_{i}^{+}|}\textstyle\sum\limits_{j\in\mathcal{G}_{i}^{+}}\frac{\exp(S_{z_{i},z_{j(i)}}/\tau)}{\mathcal{G}_{i}})$ (3)
For all negative samples in $\mathcal{G}_{i}^{-}$, we calculate the weight $w_{x_{k}}=\frac{1}{d(x_{i},x_{k})}$ for each $x_{k}\in\mathcal{G}_{i}^{-}$ based on its distance relative to anchor $x_{i}$. We then normalize these weights to obtain the probability $p_{x_{k}}$: $p_{x_{k}}=\frac{w_{x_{k}}}{\textstyle\sum_{x_{m}\in\mathcal{G}_{i}^{-}}w_{x_{m}}}$.
$\mathcal{G}_{i}=\textstyle\sum\limits_{x_{j(i)}}\exp(S_{z_{i},z_{j(i)}}/\tau)+\textstyle\sum\limits_{x_{k(i)}}p_{x_{k}}\exp(S_{z_{i},z_{k(i)}}/\tau)$ (4)
where $\tau$ is the temperature hyper-parameter and $S$ is the similarity function, typically defined by cosine similarity.
Finally, MGM and TEM-CL constitute a unified pre-training framework with a total loss:
$\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{MGM}}+\lambda\mathcal{L}_{\text{Tri}}$ (5)
where $\lambda$ is a hyper-parameter tuning the influence between two loss functions.
4 Experiments
We pre-train FGBERT on a large amount of metagenomic data and comprehensively assess its generalization on different datasets, from thousands to hundreds of thousands, which are described in detail in Table 1. Our goal is to validate the performance of our proposed model on four levels of tasks: (1) Genomic Structure Analysis; (2) Functional Gene Prediction (3) Pathogenicity Potential Assessment and (4) Nitrogen Cycle Prediction.
4.1 Pre-training Datasets
We use MGnify updated 2023_02, containing 2,973,257,435 protein sequences. MGnify [50] is a microbial genome dataset that contains genome sequences from different microbial communities. Detailed dataset information is shown in Appendix 5.
Table 1: Description of Experimental Datasets. #Seq. and #Class means the number of sequences and categories in each dataset.
Dataset
Description
#Seq.
#Class
E-K12
Gene Operons
4,315
1,379
CARD-A
AMR Gene Family
1,966
269
CARD-D
Drug Class
1,966
37
CARD-R
Resistance Mechanism
1,966
7
VFDB
Virulence Factors
8,945
15
ENZYME
Enzyme Functions
5,761
7
PATRIC
Pathogenic Genes
5,000
110
NCycDB
Nitrogen Cycling Genes
213,501
68
[…]
Method Operons ARG Prediction Virus Enzyme Pathogen N-Cycle
E-K12 CARD-A CARD-D CARD-R VFDB ENZYME PATRIC NCycDB
M.F1
W.F1
M.F1
W.F1
M.F1
W.F1
M.F1
W.F1
M.F1
W.F1
M.F1
W.F1
M.F1
W.F1
M.F1
W.F1
RF
20.2
34.8
22.4
35.3
36.1
49.0
47.8
57.6
22.4
38.5
33.6
41.2
25.3
29.8
67.0
71.7
SVM
38.6
45.2
27.6
40.5
33.6
47.2
43.3
66.2
28.0
41.4
31.3
43.6
26.6
31.2
66.9
70.3
KNN
39.9
41.0
36.9
54.4
36.4
51.3
36.2
63.5
27.3
47.1
31.4
42.9
11.0
27.4
68.8
73.2
LSTM
40.4
42.5
47.1
60.3
39.1
62.3
47.5
84.2
36.7
66.3
42.8
51.0
41.3
49.7
71.9
81.2
BiLSTM
38.2
43.8
47.4
61.9
43.5
58.1
58.9
80.3
46.1
72.1
38.7
50.2
43.3
48.5
82.0
88.4
VT
43.3
47.8
57.1
70.0
49.8
68.1
55.7
86.4
58.0
81.0
68.2
75.8
49.8
57.3
84.5
90.7
HyenaDNA
42.4
47.1
50.9
68.2
53.6
78.1
66.2
88.1
61.0
70.4
79.6
83.6
51.1
57.6
92.4
96.0
ESM-2
38.2
42.5
57.2
71.4
56.0
82.1
68.2
90.0
60.7
84.4
92.5
96.7
56.0
67.5
95.8
96.1
NT
45.1
44.8
58.5
72.0
56.2
80.2
68.0
90.3
58.3
71.6
74.1
76.7
46.1
61.9
75.1
86.5
DNABERT2
51.7
52.4
65.2
79.8
51.5
78.7
61.2
88.6
58.2
82.3
85.4
85.2
52.9
60.6
88.6
95.7
Ours 61.8 65.4 78.6 90.1 57.4 85.2 69.4 91.4 75.7 90.2 99.1 98.8 99.3 99.0 99.5 99.2
4.2 Genomic Structure Analysis
Gene Operons Prediction is to identify the transcription factor binding sites that have the strongest correlation with operon regulation in the gene regulatory network, which helps us to understand the mechanism and network of gene regulation.
Dataset. The transcription factor binding sites with the strongest correlation to the operon in E.coil K12 RegulonDB dataset (E-K12) [51] consists of 4315 operons. Detailed dataset information is listed in Appendix 6.
Results. Table 2 shows that our model achieves the SOTA results for the operon prediction task. From the attention heatmap in Figure 4, it can be observed that gene operon tolC has high attention weight with the operons ygiB and ygiC and yqiA. This suggests a significant interaction among these operons, indicating the presence of a shared genetic operon tolC-ygib. Biological research supports that these operons are indeed related to the DUF1190 domain-containing protein YgiB [28].
Figure 4: Visualization of the attention heatmap from the first head of the last (19th) attention layer. Darker blue color indicates higher attention weight.
4.3 Functional Gene Prediction
Antimicrobial Resistance Genes (ARG) Prediction is crucial for understanding antimicrobial resistance genes and facilitates the identification of resistance mechanisms. However, existing methods suffer from high false-positive rates and specific category bias [26, 5]. Therefore, there is an urgent need for deep learning methods to rapidly and accurately predict the presence of ARG in metagenomic data.
Dataset. CARD [25] is a comprehensive ARG database in which each gene is associated with a specific AMR Gene Family (CARD-A: 269 categories in total), Drug Class (CARD-D: 37 categories in total), and Resistance Mechanism(CARD-R: 7 categories in total). Therefore, our model performs a three-category classification prediction task for each gene sequence in CARD, respectively.
Results. The prediction on CARD-A category is especially noteworthy in Table 2. Because this category covers up to 269 different classifications, resulting in a long-tail issue. This requires the model to understand the biological properties of gene sequences and accurately annotate the sequences [40]. To address the long-tail problem, we adjust the data sampling strategy to increase the sampling frequency of fewer samples to improve the model’s prediction accuracy. Table 3 demonstrates the prediction accuracy of FGBERT on CARD-R category. This result shows that our model has better classification results for both majority and minority classes with more than 85% accuracy. Appendix 9 shows that FGBERT has better classification performance for all 269 categories of AMR Gene Family, with 100% data classification accuracy for majority categories such as CTX, ADC, CMY, etc. and minority categories such as AXC, CRH, KLUC, and so on.
Table 3: Detailed Classification Results of FGBERT Model on Resistance Mechanism Category Prediction, including the total number of categories, the number of correct predictions, and the accuracy rate for each class.
Class
Name
Total
Count
Correct
Num
Correct
Ratio
Antibiotic Inactivation 252 238 94.44%
Antibiotic Target Alteration 70 59 84.29%
Antibiotic Target Protection 28 27 96.43%
Antibiotic Efflux 25 18 72.00%
Antibiotic Target Replacement 14 12 85.72%
Virulence Factors (VF) Prediction is to detect microbial elements like bacterial toxins, which enhance pathogen infectivity and exacerbate antimicrobial resistance. Existing methods for analyzing metagenomic analyses are insufficient, especially tools for co-predicting ARGs and VFs are scarce, and also suffer from threshold sensitivity [61].
Dataset. VFDB [8] dataset includes the major VFs of the most characterized bacterial pathogens, as well as the structural features, functions, and mechanisms of these pathogens. We use VFDB core dataset, including 8945 VF sequences and 15 VF categories.
Results. According to the results reported in Table 2, our model achieves the SOTA results on VFDB. Compared to the genomic pre-trained model, our advantage is relatively obvious. For example, M.F1 improves by 30% and W.F1 improves by 9.5% compared to DNABERT2. This further highlights the challenge of directly transferring the genome pre-trained model to the metagenomic domain for precise functional annotation. Conversely, ESM-2, as a PLM, outperformed the genome pre-trained model, emphasizing the significance of harnessing the inherent protein information within metagenomic data.
Enzyme Function Prediction is critical for understanding metabolism and disease mechanisms in organisms. While traditional methods rely on time-consuming and labor-intensive biochemical experiments, advanced technologies can be used to efficiently and accurately predict large-scale genomic data.
Dataset. ENZYME [7] is a repository of information related to the nomenclature of enzymes. We organize 5761 data, 7 categories of ENZYME core dataset, each enzyme has a unique EC number [44].
Results. Our experimental results demonstrate FGBERT’s superior performance on ENZYME dataset. It outperforms ESM-2, the second-highest method, by approximately 6.62% in M.F1 and 2.09% in W.F1, demonstrating its ability to discern distinct enzyme function characteristics. This observation highlights that our model not only captures gene-protein contextual relationships but also effectively models the relationships between sequences and functions within metagenomic data.
4.4 Pathogenicity Potential Assessment
Genome Pathogens Prediction is to assess the pathogenic potential of pathogens to cope with the public health risks caused by newly emerging pathogens. Hence, the development of accurate deep-learning algorithms for the precise identification of pathogens can improve the ability to respond to the threat of drug resistance.
Dataset. We use PATRIC core dataset [15], which has 5000 pathogenic bacterial sequences and 110 pathogenic bacterial classes.
Results. Table 2 displays the classification results for pathogenic bacteria species on PATRIC. The result shows that FGBERT can recognize key features of genera, resulting in better classification results than baselines. PATRIC dataset, which consists of a large number of categories with relatively fewer data, poses a challenge for models. Due to this complexity, baselines underperform, as they typically require more data to discern the subtle differences between numerous categories. In contrast, FGBERT stands out with M.F1 and W.F1 scores of 99.27% and 99.03% respectively, and appears to be robust against such data scarcity and category abundance, indicating an advanced learning capability suited for high-dimensional classification tasks. This observation highlights the advantage of leveraging protein-based gene representations for improved accuracy in functional annotation tasks.
4.5 Nitrogen Cycle Prediction
Nitrogen Cycling Process Prediction is to quantitatively study the functional genes related to the N cycle and linking them to environmental and ecological processes is one of the important focuses in environmental genomics research.
Dataset. NCycDB [55] contains a total of 68 genes (sub)families and covers eight N cycle processes with 213501 representative sequences at 100% identity cutoffs, each of which involves a specific gene family and is critical for environmental and ecosystem nitrogen homeostasis.
Results. Table 2 presents the classification results of FGBERT on NCycDB. The findings suggest that FGBERT can recognize important features of a particular N cycle and improve gene family classification by recognizing domains. On NCycDB, which consists of a substantial 210,000 data from 68 categories, we observe that the baselines perform comparatively better than on PATRIC. This is due to the larger data size per category, which allows baselines to better discriminate between the diverse categories, while FGBERT still outperforms with remarkable M.F1 and W.F1 scores of 99.49% and 99.22%, respectively. However, the pre-trained baselines take more time and larger memory for the tokenization process for large datasets, which is analyzed in Sec 4.7.
4.6 Ablation Study
Figure 5: Ablation studies of proposed modules on four downstream tasks.
[…]
Method Operons ARG Prediction
E-K12 CARD-A CARD-D CARD-R
FGBERT 61.8 78.7 57.4 69.4
$w/o.$ MGM -8.1 -6.7 -10.5 -6.7
$w/o.$ Triplet -7.4 -5.4 -5.8 -3.4
Figure 6: t-SNE visualization of input protein embeddings and our proposed protein-based gene embeddings for ATP synthase sequences (n=1177). Each icon denotes an ATP synthase sequence and is colored according to different functions.
Figure 7: Comparative Analysis on Tokenization Efficiency: Balancing Time(s) vs. Memory Usage(MB). Each circle point on the plot corresponds to a specific dataset, with the point size indicating the dataset’s scale.
[…]
Furthermore, we perform a visualization experiment to validate the effectiveness of MGM in One-to-Many scenario. ATP synthases can exhibit different functions in different organisms to adapt to their respective environmental conditions, even though the basic functions are the same [22]. To explore this, we collect 1177 ATP synthase sequences from UniProt [1] and color them according to six taxonomies, i.e. Hydrolases (35.6%, dark green), Isomerases (3.6%, orange), Ligases (35.7%, blue), Lyases (10.7%, pink), Oxidoreduct (8.9%, light green) and Transferases (5.3%, yellow) (numbers in parentheses indicate category ratios). Figure 6a shows the clustering results of ATPase protein embeddings without genome contextual analysis, indicating that more dispersed clustering results of genes in different genome contexts. Figure 6b presents the clustering results of ATPase embedding after our designed protein-based gene embedding, which shows genes belonging to the same category (in the same color) apparently cluster together. Concretely, Isomerases (3.5%, orange) and Transferases (5.3%, yellow), which account for the smallest percentage, cluster together, whereas, in the left plot, these two are scattered. This demonstrates that our proposed MGM can resolve the One-to-Many relationship between sequences and functions.
4.7 Model Efficiency Study
We analyze the time complexity and memory efficiency of token processing for ours and four genome pre-trained methods on six datasets in Figure 7. Our tokenizer achieves the best balance between time, computation, and performance (M.F1) on all datasets. Notably, for NCyc dataset (brown circle) with 213,501 sequences, ours exhibits a remarkable reduction in time by 31.05% compared to DNABERT2 along with a substantial decrease in memory usage by 94.33%. Additionally, for CARD dataset (orange circle) containing 1,966 sequences, ours showcases a significant reduction in time by 61.70% compared to DNABERT2 accompanied by a substantial decrease in memory usage by 58.53%. Although HyenaDNA has a smaller memory footprint than ours on the Operons, CARD, VFDB, and ENZYME datasets, it performs worse than ours in terms of time cost, and performance.
Figure 8: Sensitivity w.r.t Hyper-parameters $\alpha,b$ of CARD dataset on AMR Gene Family.
4.8 Sensitivity Analysis
Our sensitivity analysis on the batch size $b$ indicates that the proposed model can be optimized effectively using small batches without larger performance degradation shown in Figure 8. This observation is particularly important for resource-constrained scenarios, highlighting our model maintains good performance even with limited data. We select a moderate batch size 1000 that balances computational efficiency and model performance. Next, we analyze another hyper-parameter of balance ratio $\alpha$ on CARD dataset. FGBERT obtains good performance for different values of $\alpha$, demonstrating its robustness and insensitivity to this hyper-parameter. We select $\alpha$ to be 0.4.
5 Conclusion
In this paper, we propose a new idea of protein-based gene representation, preserving essential biological characteristics within each gene sequence. With the new context-aware tokenizer, we propose MGM, a gene group-level pre-training task, designed to learn the interactions between genes. Additionally, we develop TEM-CL, a contrastive learning module to generate multiple positive and negative samples to distinguish the gene sequences. MGM and TEM-CL constitute a joint pre-training model, FGBERT for metagenomic data. Our experiments and visualizations demonstrate the superior performance of our model. For the future, it remains to be explored how to incorporate multi-omics data, such as metabolomics, into our metagenomic pre-trained model.
Appendix A Prelinaries
Transformers are the dominant tools for modeling sequence data. In Transformer, the input sequence is first transformed into a series of vectors, each corresponding to an element of the sequence (e.g., a word or character), which are generated by an embedding layer.
Subsequently, self-attention mechanism computes the attention score of each element with respect to the other elements in the sequence. Specifically, by multiplying three different weight matrices, each input vector $X$ is transformed into: a query vector $Q$, a key vector $K$, and a value vector $V$. Then, the attention score is computed.
[…]
These scores are normalized by Softmax function, where it is multiplied by the value vector $V$ to get the weighted value vector. All these weighted value vectors are summed to form the output Z of the attention layer.
[…]
Transformers typically use a multi-head attention mechanism to capture the information in the sequence more comprehensively. The outputs of each self-attention and feed-forward network layer are passed through a residual connection, followed by layer normalization. After multi-layer processing, the Transformer model outputs a series of vectors suitable for a variety of downstream tasks such as categorization, translation, or text generation of text, speech, or any form of sequence data.
Appendix B Model Configurations
All our experiments are performed on 4 NVIDIA V100 GPUs and the PyTorch framework. The encoder of FGBERT is initialized with Roberta [11]. The parameters for pre-training: batch size = 1000, 19 encoder layers, 10 attention heads, embedding dimension = 1280, and relative position encoding. During the pre-training stage, the model is trained for 500 epochs. We use the AdamW optimizer [35] with a weight decay of 0.02. The learning rate is initialized to 1e-5 and warmed up to 1e-4 after 5000 epochs. Then, it is reduced to 1e-5 by a cosine decay strategy. The overall model comprises 954.73M parameters with a computational load of 2.55B FLOPs.
Appendix C Dataset.
MGnify dataset includes genome sequence information from different microbial communities and their Biome Names (e.g., ”Engineered”, ”Host-associated”, and ”Control”), as well as their corresponding Sample Classes, Sample Numbers, and Spectrum Categories. Lineage Category refers to the taxonomic information of microorganisms, which describes the taxonomic hierarchy of microorganisms from Domain (e.g., ”root” or ”Engineered”) to Species (e.g., ”Continuous culture” or ”Saliva”), which can be observed in Table 5.
Table 5: Microbial Species and Genomic Data Across Various Environments in the MGnify Dataset. Each column is catalogs, species, and genomes respectively.
Catalogues Species Genomes
Human gut 4744 289232
Human oral 452 1225
Cow rumen 2729 5578
Marine 1496 1504
Pig gut 1376 3972
Zebrafish feacel 79 101
Non-model fish gut 172 196
E.coil K12 RegulonDB Information such as the name of the operon, its description, and the name of the genes included in the partially E. coil K12 RegulonDB dataset are listed in Table 6.
Table 6: Examples of Operons Dataset. Information such as the name of the operon, its description, and the name of the genes included in the partially E. coil K12 RegulonDB dataset
GeneID Operon name
Discription Operon name
Gene name(s)
contained in the operon
Other database’s id
related to gene
2971 ygiV
DNA-binding transcriptional repressor YgiV
None
None
2972 ygiW
BOF family protein YgiW
None
None
2973 qseB
DNA-binding transcriptional activator QseB
qseBC
None
2974 qseC
sensor histidine kinase QseC
qseBC
None
2975 ygiZ
DUF2645 domain-containing inner membrane protein YgiZ
None
None
2976 mdaB
NADPH:quinone oxidoreductase MdaB
None
None
2977 ygiN
putative quinol monooxygenase YgiN
None
None
2978 parE
DNA topoisomerase IV subunit B
nudF-yqiB-cpdA-yqiA-parE
None
2979 yqiA
esterase YqiA
nudF-yqiB-cpdA-yqiA-parE
None
2980 cpdA
cAMP phosphodiesterase
nudF-yqiB-cpdA-yqiA-parE
None
2981 yqiB
DUF1249 domain-containing protein YqiB
nudF-yqiB-cpdA-yqiA-parE
None
2982 nudF
ADP-sugar pyrophosphatase
nudF-yqiB-cpdA-yqiA-parE
None
2983 tolC
outer membrane channel TolC
tolC-ygiBC
tolC-ygiBC
2984 ygiB
DUF1190 domain-containing protein YgiB
tolC-ygiBC
tolC-ygiBC
2985 ygiC
putative acid–amine ligase YgiC
tolC-ygiBC
tolC-ygiBC
2986 ygiD
4,5-DOPA dioxygenase extradiol
None
None
  • Note: This table provides a partial list of instances from the Operons dataset, including Gene data IDs, operon names, descriptions, Gene name(s) contained in the operon, and other database’s id related to the gene.
CARD Table 7 shows a partial list of instances from the CARD dataset, including data IDs, ARO (Antibiotic Resistance Ontology) Accession IDs, AMR (Antimicrobial Resistance) category names, drug category names, and Resistance Mechanism names.
Table 7: Examples of CARD Dataset
Model ID ARO Accession
AMR Gene Family
Drug Class
Resistance Mechanism
2 3002999
CblA beta-lactamase
cephalosporin
antibiotic inactivation
5 3002867
trimethoprim resistant dihydrofolate reductase dfr
diaminopyrimidine antibiotic
antibiotic target replacement
7 3001989
CTX-M beta-lactamase
cephalosporin
antibiotic inactivation
10 3002244
CARB beta-lactamase
penam
antibiotic inactivation
20 3002012
CMY beta-lactamase
cephamycin
antibiotic inactivation
24 3003552
Target protecting FusB-type protein conferring resistance to Fusidic acid
fusidane antibiotic
antibiotic target protection
25 3001980
CTX-M beta-lactamase
cephalosporin
antibiotic inactivation
27 3002835
lincosamide nucleotidyltransferase (LNU)
lincosamide antibiotic
antibiotic inactivation
31 3003163
CTX-M beta-lactamase
cephalosporin
antibiotic inactivation
35 3002804
fosfomycin thiol transferase
phosphonic acid antibiotic
antibiotic inactivation
38 3002649
APH(3’)
aminoglycoside antibiotic
antibiotic inactivation
39 3002528
AAC(3)
aminoglycoside antibiotic
antibiotic inactivation
40 3002771
quinolone resistance protein (qnr)
fluoroquinolone antibiotic
antibiotic target protection
41 3003198
16S rRNA methyltransferase (G1405)
aminoglycoside antibiotic
antibiotic target alteration
46 3002125
CMY beta-lactamase
cephamycin
antibiotic inactivation
50 3002380
SME beta-lactamase
carbapenem
antibiotic inactivation
51 3002538
AAC(3)
aminoglycoside antibiotic
antibiotic inactivation
58 3002762
quinolone resistance protein (qnr)
fluoroquinolone antibiotic
antibiotic target protection
60 3002795
quinolone resistance protein (qnr)
fluoroquinolone antibiotic
antibiotic target protection
63 3002546
AAC(6’)
aminoglycoside antibiotic
antibiotic inactivation
64 3002083
CMY beta-lactamase
cephamycin
antibiotic inactivation
69 3002620
ANT(3”)
aminoglycoside antibiotic
antibiotic inactivation
71 3002778
quinolone resistance protein (qnr)
fluoroquinolone antibiotic
antibiotic target protection
75 3003026
fusidic acid inactivation enzyme
fusidane antibiotic
antibiotic inactivation
88 3002066
CMY beta-lactamase
cephamycin
antibiotic inactivation
92 3001904
CTX-M beta-lactamase
cephalosporin
antibiotic inactivation
94 3002092
CMY beta-lactamase
cephamycin
antibiotic inactivation
95 3002067
CMY beta-lactamase
cephamycin
antibiotic inactivation
98 3002059
CMY beta-lactamase
cephamycin
antibiotic inactivation
99 3002753
quinolone resistance protein (qnr)
fluoroquinolone antibiotic
antibiotic target protection
105 3002243
CARB beta-lactamase
penam
antibiotic inactivation
106 3002681
chloramphenicol acetyltransferase (CAT)
phenicol antibiotic
antibiotic inactivation
110 3002565
AAC(6’)
aminoglycoside antibiotic
antibiotic inactivation
114 3003012
trimethoprim resistant dihydrofolate reductase dfr
diaminopyrimidine antibiotic
antibiotic target replacement
116 3002733
quinolone resistance protein (qnr)
fluoroquinolone antibiotic
antibiotic target protection
120 3002604
ANT(3”)
aminoglycoside antibiotic
antibiotic inactivation
124 3002263
IND beta-lactamase
carbapenem
antibiotic inactivation
128 3002828
Miscellaneous ABC-F subfamily ATP-binding cassette ribosomal protection proteins
macrolide antibiotic
antibiotic target protection
129 3000198
fosfomycin thiol transferase
phosphonic acid antibiotic
antibiotic inactivation
130 3001972
CTX-M beta-lactamase
cephalosporin
antibiotic inactivation
  • Note: This table provides a partial list of instances from the CARD dataset, including data IDs, ARO (Antibiotic Resistance Ontology) Accession IDs, AMR (Antimicrobial Resistance) category names, drug category names, and Resistance Mechanism names.
Appendix D Baselines
We evaluate ours compared with three traditional machine learning methods (RandomForest, SVM, KNN), three sequence modeling methods (BiLSTM [52] vs. LSTM [21]), Vanilla Transformer and four pre-trained models on protein/DNA sequences(ESM-2 [33], HyenaDNA [43], NT [10] and DNABERT2 [63]) for comparison. Machine learning methods encode sequences by K-mers, where K takes the best one of the [3,4,5,6] parameters. Sequence modeling methods encode genes using Word2vec. For a fair comparison, each model was trained and optimized to obtain the corresponding optimal hyperparameters. We adopt a 5-fold cross-validation approach to partition all downstream task datasets into a training set (80%), a validation set (10%), and a test set (10%).
Appendix E Evaluation Metrics.
To fully evaluate the model performance, we used several evaluation metrics including Accuracy, Macro Pre, Macro Recall, Macro F1 and Weighted F1.
Precision: indicates the precision rate, which is the ratio of the number of true positive samples to the number of samples classified as positive. precision = TP / (TP + FP)
Recall: the ratio of the number of true positive samples to the number of all positive samples in the sample set, Recall = TP / (TP + FN)
F1score: used to combine Precision and Recall, F1-Score = 2 * Precision * Recall / (Precision + Recall)
Appendix F Overview of Downstream Tasks
Operon Prediction Task is to identify the transcription factor binding sites that have the strongest correlation with operon regulation in the gene regulatory network [9, 13], which helps us to understand the mechanism and network of gene regulation and reveals the key interactions between transcription factors and operons [46]. In general, operons are composed of a set of multiple genes close to each other.
ARGs and VFs Prediction Task. Pathogenic microorganisms pose a threat to public health by invading hosts through virulence factors (VFs) and exacerbating antibiotic resistance genes (ARGs) [45]. Despite different evolutionary pathways, VFs and ARGs share common features that are critical for pathogenic bacteria to adapt to and survive in the microbial environment [38]. Therefore, accurate identification of VFs and ARGs is extremely important for understanding the relationship between the microbiome and disease. However, traditional sequence-matching-based ARG identification methods suffer from high false-positive rates and specific ARG category bias [26, 5]. Existing methods for analyzing metagenomic data are insufficient, especially tools for co-predicting the two are scarce, and also suffer from threshold sensitivity [61]. Therefore, there is an urgent need for deep learning methods to rapidly and accurately predict the presence of VF and ARG in metagenomic datasets.
Overview of Enzymes. Enzymes are important catalysts in living cells that produce essential molecules needed by living organisms through chemical reactions [54]. While traditional methods rely on time-consuming and labor-intensive biochemical experiments, advanced technologies such as deep learning can be used to efficiently and accurately predict large-scale genomic data, accelerating biomedical research and new drug development.
Overview of Pathogens. Pathogenic bacteria are a group of bacteria that can cause disease and can pose a significant threat to human health compared to commensal non-pathogenic bacteria in the human body with their ability to invade the host and cause disease [27]. The existing methods are limited to long genome sequences and are dominated by short sequence (¡3 kb) inputs, which are unable to directly process long genome sequences at the million bp level. Hence, the development of accurate deep learning algorithms for the precise identification of pathogens (causative agents) can help in the development of new therapies and vaccines, and improve the ability to respond to the threat of drug resistance.
Overview. The Nitrogen cycle is a collection of important biogeochemical pathways in the Earth’s ecosystems, and quantitatively studying the functional genes related to the N cycle and linking them to environmental and ecological processes is one of the important focuses in environmental genomics research [16]. Currently, macrogenome sequencing has been widely used to characterize gene families involved in N cycling processes. However, existing methods usually face the problems of inefficient database searches, unclear direct lineage classification, and low coverage of N cycle genes and/or gene (sub)families when analyzing N cycle gene families in macrogenomes. Consequently, accurate prediction of the functional categories of genes or proteins related to nitrogen metabolism in microorganisms is important for analyzing important resources in the process of nitrogen cycling and understanding the nitrogen cycle in ecosystems.
Appendix G Detailed Classification Results on CARD Dataset
Table 8: Detailed Classification Results of FGBERT Model on Drug Class Category Prediction.
Class Name Total Count Correctly Num Correct Ratio
cephalosporin 103 101 98.06%
aminoglycoside antibiotic 47 37 78.72%
cephamycin 34 31 91.18%
fluoroquinolone antibiotic 33 27 81.82%
carbapenem 31 25 80.65%
peptide antibiotic 30 26 86.67%
penam 23 21 91.30%
tetracycline antibiotic 16 10 62.50%
macrolide antibiotic 12 10 83.33%
diaminopyrimidine antibiotic 11 8 72.73%
phenicol antibiotic 11 6 54.55%
phosphonic acid antibiotic 10 6 60.00%
lincosamide antibiotic 3 0 0.00
rifamycin antibiotic 3 1 33.33%
disinfecting agents antiseptics 3 2 66.67%
nitroimidazole antibiotic 3 1 33.33%
aminocoumarin antibiotic 2 0 0.00%
sulfonamide antibiotic 2 0 0.00%
elfamycin antibiotic 2 2 100.00%
isoniazid-like antibiotic 2 1 50.00%
fusidane antibiotic 1 0 0.00%
nucleoside antibiotic 1 0 0.00%
polyamine antibiotic 1 0 0.00%
thioamide antibiotic 1 0 0.00%
glycopeptide antibiotic 1 0 0.00%
mupirocin-like antibiotic 1 0 0.00%
pleuromutilin antibiotic 1 0 0.00%
salicylic acid antibiotic 1 0 0.00%
Table 9: Detailed Classification Results of FGBERT Model on AMR Gene Family Category Prediction. Our model performs a three-class classification prediction task for each gene sequence in the CARD dataset, covering AMR Gene Family (269 classes in total), Drug Class (37 classes in total) and Resistance Mechanism (7 classes in total). In order to comprehensively evaluate the model performance, we used several evaluation metrics, including Accuracy, Macro Pre, Macro Recall, Macro F1, Weighted Pre, Weighted Recall and Weighted F1. One of the challenging tasks is to predict the AMR Gene Family category, which has up to 269 different classes and requires high complexity classification of gene sequences with multiple labels. This requires the model to be able to understand the biological features of the gene sequences and accurately map them to the corresponding AMR Gene Family categories to ensure accurate prediction in the multi-classification task.
Class Name Total Count Correct Num Correct Ratio
CTX-M beta-lactamase 46 46 100.00%
ADC beta-lactamases pending classification for carbapenemase activity 36 36 100.00%
CMY beta-lactamase 33 33 100.00%
quinolone resistance protein (qnr) 22 22 100.00%
MCR phosphoethanolamine transferase 21 20 95.24%
major facilitator superfamily (MFS) antibiotic efflux pump 17 15 88.24%
trimethoprim resistant dihydrofolate reductase dfr 12 8 66.67%
AAC(6’) 12 12 100.00%
CARB beta-lactamase 10 9 90.00%
ADC beta-lactamase without carbapenemase activity 10 3 30.00%
16s rRNA with mutation conferring resistance to aminoglycoside antibiotics 9 7 77.78%
fosfomycin thiol transferase 6 3 50.00%
ANT(3”) 6 6 100.00%
chloramphenicol acetyltransferase (CAT) 6 5 83.33%
23S rRNA with mutation conferring resistance to macrolide antibiotics 5 5 100.00%
APH(3’) 4 2 50.00%
AAC(3) 4 4 100.00%
IND beta-lactamase 4 4 100.00%
ATP-binding cassette (ABC) antibiotic efflux pump 4 3 75.00%
tetracycline-resistant ribosomal protection protein 4 3 75.00%
IMI beta-lactamase 4 4 100.00%
CfiA beta-lactamase 4 4 100.00%
fluoroquinolone resistant gyrA 4 2 50.00%
16S rRNA methyltransferase (G1405) 3 1 33.33%
resistance-nodulation-cell division(RND) antibiotic efflux pump 3 0 0.00%
macrolide phosphotransferase (MPH) 3 0 0.00%
tetracycline inactivation enzyme 3 3 100.00%
fluoroquinolone resistant parC 3 3 100.00%
lincosamide nucleotidyltransferase (LNU) 2 1 50.00%
rifampin ADP-ribosyltransferase (Arr) 2 2 100.00%
CphA beta-lactamase 2 2 100.00%
AAC(2’) 2 2 100.00%
FRI beta-lactamase 2 2 100.00%
HERA beta-lactamase 2 2 100.00%
EC beta-lactamase 2 2 100.00%
nitroimidazole reductase 2 1 50.00%
Target protecting FusB-type protein conferring resistance to Fusidic acid 1 1 100.00%
SME beta-lactamase 1 1 100.00%
Miscellaneous ABC-F subfamily ATP-binding cassette ribosomal protection proteins 1 0 0.00%
APH(2”) 1 0 0.00%
CfxA beta-lactamase 1 1 100.00%
SRT beta-lactamase 1 1 100.00%
aminoglycoside bifunctional resistance protein 1 0 0.00%
APH(9) 1 0 0.00%
sulfonamide resistant sul 1 0 0.00%
ANT(6) 1 0 0.00%
multidrug and toxic compound extrusion (MATE) transporter 1 0 0.00%
methicillin resistant PBP2 1 1 100.00%
streptothricin acetyltransferase (SAT) 1 0 0.00%
pmr phosphoethanolamine transferase 1 1 100.00%
BlaZ beta-lactamase 1 0 0.00%
rifampin monooxygenase 1 0 0.00%
APH(6) 1 0 0.00%
macrolide esterase 1 0 0.00%
ANT(4’) 1 0 0.00%
ANT(9) 1 0 0.00%
AQU beta-lactamase 1 1 100.00%
CepA beta-lactamase 1 1 100.00%
APH(3”) 1 0 0.00%
sulfonamide resistant dihydropteroate synthase folP 1 0 0.00%
aminocoumarin resistant gyrB 1 0 0.00%
fluoroquinolone resistant parE 1 0 0.00%
daptomycin resistant cls 1 0 0.00%
16s rRNA with mutation conferring resistance to peptide antibiotics 1 0 0.00%
elfamycin resistant EF-Tu 1 0 0.00%
defensin resistant mprF 1 1 100.00%
antibiotic resistant ndh 1 0 0.00%
fluoroquinolone resistant gyrB 1 0 0.00%
16S rRNA with mutation conferring resistance to tetracycline derivatives 1 0 0.00%
FONA beta-lactamase 1 1 100.00%
antibiotic-resistant isoleucyl-tRNA synthetase (ileS) 1 1 100.00%
antibiotic-resistant murA transferase 1 0 0.00%
antibiotic-resistant UhpT 1 1 100.00%
small multidrug resistance (SMR)antibiotic efflux pump 1 0 0.00%
ARL Beta-lactamase 1 1 100.00%
General Bacterial Porin with reduced permeability to peptide antibiotics 1 0 0.00%
CMH beta-lactamase 1 1 100.00%
AXC beta-lactamase 1 1 100.00%
CRH beta-lactamase 1 1 100.00%
KLUC beta-lactamase 1 1 100.00%
LUT beta-lactamase 1 1 100.00%
PFM beta-lactamase 1 1 100.00%
SGM beta-lactamase 1 1 100.00%